OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2 pi] -- 2cos(x) + sin(2x) = 0 ?? Please HELP!

jimthompson5910 (jim_thompson5910):

use the identity sin(2x) = 2*sin(x)*cos(x)

jimthompson5910 (jim_thompson5910):

so 2cos(x) + sin(2x) = 0 turns into 2cos(x) + 2*sin(x)*cos(x) = 0

OpenStudy (anonymous):

thankyou (:

jimthompson5910 (jim_thompson5910):

then you factor out 2cos(x) to get 2cos(x) * (1 + sin(x)) = 0

jimthompson5910 (jim_thompson5910):

np